2x+2y=7
x+y=6
solve the following systems using substitution or elimination

Answer by elima(1433)

2x+2y=7
x+y=6
first lets solve the second equation for x;
x+y=6
x=-y+6
Now substitute that into the first equation for x;
2(-y+6)+2y=7
-2y+6+2y=7
-2y+2y=1
no solution
:)